OpenStudy (anonymous):

A box with an open top is to be constructed from a square piece of cardboard, 3 ft wide, by cutting out a square from each of the four corners and bending up the sides. Find the largest volume that such a box can have. Let x denote the length of the side of the square being cut out. Let y denote the length of the base. write the expression for the volume V: write the volume as a function of x: finish solving the problem by finding the largest volume the box can have:

OpenStudy (anonymous):

I saw some answers online but I didn't understand their responses, I was wondering if someone could explain it to me a bit more clearly :)

OpenStudy (anonymous):

I get that it would require an actual explanation but this is kind of the best I can do at the moment?

hero (hero):

Okay, I will try, but it might take 5-10 minutes

OpenStudy (anonymous):

okay thank you

hero (hero):

Are you familiar with taking derivatives?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I remember doing this last year when I was in calculus I just don't remember the steps exactly because it's been so long, you don't have to go too in depth

OpenStudy (anonymous):

for example, I think that the volume should be : (3-2x)(3-2y)(x) but I'm trying to enter it into my online homework and it isn't working and I'm not sure how to continue from there

hero (hero):

Yes, that's not right because you only need to deal with one variable: x

hero (hero):

Remember, we're dealing with a square, not a rectangle

hero (hero):

so V = x(3-2x)^2

OpenStudy (anonymous):

then I would get \[9x-12x^2+4x^3\] i tried that too and it won't work for the part that says expression of the volume

hero (hero):

Okay, there are other forms to can write for the expression of V: \[V = x(3-2x)(3-2x)\] \[V = x(3-2x)^2\] Try those

OpenStudy (anonymous):

yeahh haha uhm none of them work

OpenStudy (anonymous):

and I only have two submissions left

hero (hero):

I'm sorry to hear that. I don't know what they want, but the expressions are not incorrect

OpenStudy (anonymous):

yeah I didn't think they were wrong either. could you just show me how to write it as a function of x and how I would eventually get what the volume is?

hero (hero):

\[V(x) = x(3-2x)^2\]

hero (hero):

Try that

OpenStudy (anonymous):

mm no that doesn't work

hero (hero):

Well, you need to just skip that problem because it only wants one input that no one will be able to figure out

OpenStudy (anonymous):

well I mean that is the right beginning, and there's the other parts that still have to be answered

OpenStudy (anonymous):

its just not being taken as the right answer for some reason

OpenStudy (anonymous):

but I still want to know how to relate to writing it as a function of x and how I would get to the final answer

hero (hero):

I can't help you with this if that's the kind of problem you're having. Obviously, the program doesn't like accepting right answers.

OpenStudy (anonymous):

haha okay thanks anyway..

OpenStudy (anonymous):

maya.y I hope that the attached solution is what you are looking for.

hero (hero):

I don't see anywhere where it says the units has to be in inches

hero (hero):

So I don't understand your approach in that regard

OpenStudy (anonymous):

Well they can convert to feet if they want to.

hero (hero):

i'm just saying...why would you give a solution in terms of inches? What led you to that approach?

OpenStudy (anonymous):

Avoids fractions.

OpenStudy (anonymous):

Hero, What do you think the presentation?

hero (hero):

It was pretty good but only if you consider your approach. I think it would have been more complete if you converted back to feet at the end

OpenStudy (anonymous):

Thanks for the response.